When a honeybee flies through the air, it develops a charge of +15pC. How many electrons did it lose in the process of acquiring this charge?
Answer #1

Solition)

Given,

Charge, q=+15 pC = 15*10^-12 C

We know, Charge, q= n*e

Where, e= charge on electron=1.6*10^-19 C

So, Number of electrons, n= q/e

=(15*10^-12 C)/(1.6*10^-19 C)= 9.37*10^7 electrons (Ans)